OpenStudy (laddiusmaximus):

y= x^2 + 4x +3/ sqrt x

OpenStudy (anonymous):

What do you need to do with that?

OpenStudy (laddiusmaximus):

oh sorry differentiate the function

OpenStudy (anonymous):

Ok, do you remember the differentiation rules?

OpenStudy (laddiusmaximus):

Yeah I know the quotient rule.

OpenStudy (anonymous):

Great. As that is a sum (the easier case), you can split it intro 3 parts and differentiate them separately.

OpenStudy (laddiusmaximus):

I got to here and was stuck. I was shown how to get the answer, but I dont understand how they did part of it.

OpenStudy (laddiusmaximus):

heres what I have, (sqrtx)(2x+4)-(x^2+4+3) 1/2x^-1/2

OpenStudy (laddiusmaximus):

over (sqrtx)^2

OpenStudy (laddiusmaximus):

those square roots are what is confusing me I dont know what to do with them.

OpenStudy (laddiusmaximus):

i changed the one on the far right hand side to 1/2 x^-1/2 so I could use the derivative

OpenStudy (anonymous):

To make it easier you can think of them as (something)^1/2. Then you can differentiatie it as if it was a power.

OpenStudy (laddiusmaximus):

right. but after I do that im lost

OpenStudy (anonymous):

When you have to differentiate\[3. (x)^{\frac{-1}{2}}\]?

OpenStudy (anonymous):

That is where you get lost?

OpenStudy (laddiusmaximus):

well yes. because that last square root you have to take the derivative. but after that, I dont know what to do. or at least the explanation in the book doesnt make any sense.

OpenStudy (laddiusmaximus):

i wish i could take a picture of the answer the book gives and show you so you can explain it

OpenStudy (anonymous):

It would be:\[3 . \frac{-1}{2} . (x)^\frac{-3}{2} . x \prime\] Which you can rewrite as \[\frac{-3}{2} . \frac{ 1 }{ \sqrt[3]{x^2}}\]

OpenStudy (laddiusmaximus):

ok can you write this whole thing out? i apologize but I need to see the whole thing worked through to get it.

OpenStudy (anonymous):

Sure. \[y = x^2 + 4x + \frac{3}{\sqrt{x}}\] So \[y \prime = (x^2)\prime + (4x)\prime + (\frac{3}{\sqrt{x}})\prime\] Because it is a sum, right?

OpenStudy (laddiusmaximus):

before you get to far into this, you know there is a sqrt x in the denominator, right?

OpenStudy (anonymous):

Oh, like this? \[y = \frac{x^2 + 4x + 3}{\sqrt{x}}\]

OpenStudy (laddiusmaximus):

correct. I apply the quotient rule, and then I get stuck

OpenStudy (anonymous):

\[y \prime = \frac{(x^2+4x+3)\prime . \sqrt{x} - (x^2+4x+3) . \sqrt{x}\prime }{\sqrt{x}^2}\]

OpenStudy (laddiusmaximus):

good so far. after this is when I get wonky \

OpenStudy (anonymous):

So far you're good, right?

OpenStudy (laddiusmaximus):

correct

OpenStudy (anonymous):

Ok, let's take it by pieces so it's easier for you to see.

OpenStudy (anonymous):

The sum is a nice one to differenciate, because you can split it into even smaller parts. \[(x^2+3x+3)\prime = 2x + 3 + 0\]

OpenStudy (anonymous):

You're good with that one?

OpenStudy (laddiusmaximus):

yup

OpenStudy (laddiusmaximus):

you mean 4x right?

OpenStudy (anonymous):

Cool. Now let's go with the nasty square root.

OpenStudy (anonymous):

Ups, yeah. My bad there!

OpenStudy (laddiusmaximus):

no worries please continue

OpenStudy (anonymous):

Ok, so the little trick with the square roots, or any root, is to think of them as a power, right?

OpenStudy (anonymous):

Because you already have a rule to differentiate X to some power, and it's relatively easy...

OpenStudy (laddiusmaximus):

x^1/2 right

OpenStudy (anonymous):

Right. So applying the same method you get: \[y \prime = (x^\frac{-1}{2})\prime = \frac{-1}{2} . X^{\frac{1}{2} - 1}\]

OpenStudy (laddiusmaximus):

1/2x^-1/2 right

OpenStudy (anonymous):

That is the same: \[y \prime = \frac{-1}{2} X^{\frac{-3}{2}}\]

OpenStudy (anonymous):

Yep, there's a minus missing.

OpenStudy (anonymous):

I'm awful with this editor.

OpenStudy (laddiusmaximus):

wait why is it -1/2?

OpenStudy (laddiusmaximus):

if sqrtx is x^-1/2, how does it become -1/2x^-1/2?

OpenStudy (anonymous):

I made a mistake again. xD You're totally right

OpenStudy (anonymous):

It's a positive 1/2

OpenStudy (laddiusmaximus):

well at least im paying attention. so its 1/2x^-1/2 right?

OpenStudy (anonymous):

\[y \prime = \frac{1}{2}X^{\frac{-1}{2}}\]

OpenStudy (laddiusmaximus):

ok. after this is where im stuck

OpenStudy (anonymous):

Now you can leave as it is, or you can change it back into a square root. And then plug it into the big quotient rule.

OpenStudy (anonymous):

\[y \prime = \frac{1}{2} \frac{1}{\sqrt{x}} = \frac{1}{2 \sqrt{x}}\]

OpenStudy (anonymous):

That's all the differentiation done. You just have to plug it in and make it look nice. :P

OpenStudy (laddiusmaximus):

so its sqrtx(2x+4)-(x^2+4x+3)-1/2sqrtx ?

OpenStudy (laddiusmaximus):

over sqrtx^2?

OpenStudy (anonymous):

\[y \prime = \frac{(2x+4) . \sqrt{x} - (x^2 + 4x + 3) . \frac{1}{2 \sqrt{x}}} {x}\]

OpenStudy (anonymous):

Exactly

OpenStudy (laddiusmaximus):

wait so the sqrtx and the sqrtx^2 cancel?

OpenStudy (anonymous):

\[y \prime = \frac{(2x+4)\sqrt{x} - {(x^2 + 4x + 3)}}{x 2\sqrt{x}}\]

OpenStudy (anonymous):

Yes. Actually, you should put a module there.

OpenStudy (laddiusmaximus):

im still lost. where did the numerator 1/2sqrtx go?

OpenStudy (anonymous):

\[y \prime = \frac{(2x+4)\sqrt{x} - (x^2 + 4x +3)}{|x| 2\sqrt{x}}\]

OpenStudy (anonymous):

To the denominator.

OpenStudy (laddiusmaximus):

so they canceled?

OpenStudy (anonymous):

No, it's stil there.

OpenStudy (anonymous):

When you multiply by 1 over something, it is the same as dividing by that something.

OpenStudy (anonymous):

So you can just put it in the denominator and it won't affect the result.

OpenStudy (laddiusmaximus):

oh right

OpenStudy (anonymous):

I hope I didn't confuse you more than you were. With my typos and all...

OpenStudy (laddiusmaximus):

so what happened to the sqrtx^2 in the denominator?

OpenStudy (anonymous):

You could leave it like that if you want to. But also you can simplify it by writing it as module of X.

OpenStudy (anonymous):

If you take any number, square it, and then take the square root it is the same as if you just grabbed that number and made it always positive.

OpenStudy (anonymous):

If it's easier for you, you can always leave it as the square root squared. It's not wrong.

OpenStudy (laddiusmaximus):

ok. what does it look like now?

OpenStudy (anonymous):

Check above, the last formula. That would be the final result.

OpenStudy (anonymous):

Do you get it or you still have doubts?

OpenStudy (laddiusmaximus):

well its not the right answer so yeah I still am confused. dont worry about it. im just going to ask the professor monday.

OpenStudy (anonymous):

What does the answer say?

OpenStudy (anonymous):

Is it much different?

OpenStudy (laddiusmaximus):

3x^2+4x-3/2x^3/2

OpenStudy (anonymous):

Oh, that's probably the result after distributing the square root of X in the numerator and then doing the division.

OpenStudy (anonymous):

In the denominator if you multiply the 2 square root of X squared by X you get the 2 . (x)^3/2

OpenStudy (laddiusmaximus):

yeah im still lost. im just going to move on.

OpenStudy (anonymous):

Yeah, take a look at it later. But it's probably a different notation, if you got the quotient rule then you could do it just fine.

OpenStudy (laddiusmaximus):

im not so sure. my algebra is pretty weak.

OpenStudy (anonymous):

Don't worry, you'll get it. It's just a matter of practice. Then it becomes pretty mechanical.

OpenStudy (laddiusmaximus):

well this is my third time taking calculus and its still hard as hell. maybe its the adhd.

OpenStudy (anonymous):

I totally get it, I took it twice, too. But perseverance is everything in this stuff. Don't give up.

OpenStudy (laddiusmaximus):

wasnt planning on it. just frustrated

OpenStudy (laddiusmaximus):

getting tired of studying for hours on end and getting crappy grades

OpenStudy (anonymous):

Right. It sucks.

OpenStudy (anonymous):

But sometimes it comes a time where you magically start understanding.

OpenStudy (laddiusmaximus):

well anytime would be great. im wasting money here.

OpenStudy (anonymous):

And you are like "Wow! I can't believe it was that easy"